# Learning Enhanced Ensemble Filters

### v1

@@ 5.1 Experimental Set-Up
In the following three Subsections 5.2, 5.4, and 5.3, we compare our method against several benchmark approaches:
@@
    EnKF Perturbed Observation (EnKF) [8, 9]: This is the classic implementation of EnKF that employs a stochastic update approach, where the observations are perturbed by adding random noise samples drawn from the observation error distribution.
@@
    Ensemble Square Root Filter (ESRF) [47]: This method is a deterministic variant of the EnKF. It directly transforms the ensemble using matrix square root operations to achieve the match the second order statistics of the EnKF, but without requiring stochastic perturbations.
@@
    Local Ensemble Transform Kalman Filter (LETKF) [14]: This method builds upon the ESRF by incorporating spatial localization. We employ the Gaspari–Cohn (GC) function [45] $\mathrm{GC}(d;\sqrt{10/3}r)$ to map distances to weights, where $r$ is the localization radius and $\sqrt{10/3}$ is a normalization constant.
@@
    Pretrained MNMEF: Trained once at $N=10$ for $1000$ epochs and applied directly to all ensemble sizes, demonstrating transfer capability across different ensemble configurations.
@@
    Fine-tuned MNMEF: Fine tune the pretrained model specific to each target ensemble size for $20$ epochs, as detailed in Subsection 4.4.
For fair comparison, each benchmark method is configured with optimized hyperparameters, which were determined through comprehensive grid search experiments conducted at various ensemble sizes. The detailed results of these grid search experiments, including parameter spaces explored and grid search heatmaps, are presented in B.3. This approach ensures that all benchmark methods achieve close to their optimal performance; the benchmarks thus present a robust test of our proposed methodology. Notably, we performed hyperparameter selection for benchmarks directly on test set performance without using a validation set, which only strengthens the benchmark methods’ advantage in our comparisons. All benchmark results were generated using the DAPPER package [48].
The performance is evaluated by the difference between the estimated state (i.e., the mean of particles) and the ground truth state over the entire trajectory. In the continuous-time setting, let $\mathbf{v}^{\dagger}(t):[0,T]\rightarrow\mathbb{R}^{d_{v}}$ denote the ground truth trajectory between time 00 and $T$, and let $\mathbf{v}(t):[0,T]\rightarrow\mathbb{R}^{d_{v}}$ denote the estimated state trajectory. Let $\|\cdot\|_{2}$ denote the Euclidean norm in $\mathbb{R}^{d_{v}}$. Then, the performance can be evaluated using the relative $L^{1}:=L^{1}([0,T];\mathbb{R}^{d_{v}})$ error:
@@
In the discrete-time setting, we interpret the discrete time index $j$ as time $t_{j}=j\Delta t$, where $\Delta t=T/J$ is the time step size. The estimated trajectory is given by the sequence $V^{(N)}=\{\bar{v}^{(N)}_{j}\}_{j=1}^{J}$ where $\bar{v}^{(N)}_{j}=\frac{1}{N}\sum_{n=1}^{N}v^{(n)}_{j}$, and the ground truth trajectory by $V^{\dagger}=\{v_{j}^{\dagger}\}_{j=1}^{J}$. The relative discrete $L^{1}$ error can then be defined as
$\mathcal{E}(V^{(N)},V^{\dagger})=\frac{\sum_{j=1}^{J}\|\mathbf{v}(j\Delta t)-\mathbf{v}^{\dagger}(j\Delta t)\|_{2}\,\Delta t}{\sum_{j=1}^{J}\|\mathbf{v}^{\dagger}(j\Delta t)\|_{2}\,\Delta t}=\frac{\sum_{j=1}^{J}\|\bar{v}_{j}-v_{j}^{\dagger}\|_{2}}{\sum_{j=1}^{J}\|v_{j}^{\dagger}\|_{2}}.$ (5.2)
This error metric can be interpreted as a form of relative root-mean-square error (R-RMSE). At each time step $j$, the quantity $\|\bar{v}_{j}-v_{j}^{\dagger}\|_{2}$ corresponds to the root-mean-square error over the $d_{v}$-dimensional state vector. The overall metric thus represents a time-averaged spatial RMSE, normalized by the magnitude of the ground truth trajectory. In what follows, we refer to this error $\mathcal{E}(V,V^{\dagger})$ as the relative RMSE (R-RMSE).
For $M_{\mathrm{test}}$ trajectories $\{V_{m}^{\dagger}\}_{m=1}^{M_{\mathrm{test}}}$ and their corresponding estimated trajectories $\{V_{m}^{(N)}\}_{m=1}^{M_{\mathrm{test}}}$, the averaged R-RMSE $\bar{\mathcal{E}}$ is given by:
¯$\bar{\mathcal{E}}=\frac{1}{M_{\mathrm{test}}}\sum_{m=1}^{M_{\mathrm{test}}}\mathcal{E}(V_{m}^{(N)},V_{m}^{\dagger}).$ (5.3)
In the experiments presented in the remainder of this section, we adopt the averaged R-RMSE $\bar{\mathcal{E}}$ as the primary evaluation metric, allowing us to ameliorate the effect of randomness inherent in each individual run and providing a more reliable assessment of the overall performance of different methods. Unless otherwise specified, the number of test trajectories is fixed at $M_{\mathrm{test}}=64$.
To further highlight the advantages of our method, we consider the relative improvement in terms of R-RMSE $\mathcal{E}$ defined as:
$\mathcal{E}_{\mathrm{RI}}=\frac{\bar{\mathcal{E}}_{\mathrm{benchmark}}-\bar{\mathcal{E}}_{\mathrm{ours}}}{\bar{\mathcal{E}}_{\mathrm{benchmark}}}.$ (5.4)
@@
We apply the following general hyperparameter setting in all of our experiments unless otherwise specified. The dynamic operator $\Psi$ in Equation (2.1) is obtained by solving a differential equation using the fourth-order Runge–Kutta (RK4) [49] method. The time step $\Delta t$ denotes the interval between consecutive observations. For numerical integration, we use a finer step size within $\Delta t$ to run the RK4 algorithm more accurately. The dynamic noise $\xi\sim\mathcal{N}(0,\Sigma)$ is typically set at $\Sigma=\sigma_{v}^{2}I$ where $I$ is the identity matrix and $\sigma_{v}=10^{-3}$ or $0$ For the observation noise, $\eta\sim\mathcal{N}(0,\Gamma)$, we define its covariance matrix as $\Gamma=\sigma_{y}^{2}I$, where we set $\sigma_{y}$ to either $0.7$ or $1.0$.
@@ 5.2 Lorenz ’96
$\frac{dv_{i}}{dt}=\bigr{(}v_{i+1}-v_{i-2}\bigl{)}v_{i-1}-v_{i}+F,\quad i=1,2,\ldots,d.$ (5.5)
@@
States $v=(d_{1},d_{2},\ldots,d_{40})\in\mathbb{R}^{d},\quad d=40$
Observations $h(v)=(v_{4},v_{8},\cdots,v_{36},v_{40})\in\mathbb{R}^{d_{\mathrm{obs}}},\quad d_{\mathrm{obs}}=10$
Obs Time Step $\Delta t=1$
(a) $\sigma_{y}=0.7$
(b) $\sigma_{y}=1$
(c) Relative improvement
@@
Figure 1: Comparison results on the Lorenz ’96 system. Panels (a) and (b) show direct RMSE comparisons between different methods at observation noise levels $\sigma_{y}=0.7$ and $\sigma_{y}=1$, respectively. Panel (c) illustrates the relative improvement of our fine-tuning method compared to benchmarks. Our proposed MNMEF method is highlighted in the legends with bold font and an asterisk (e.g. Pretrain).
@@
In Figure 1 we compare the performance, using R-RMSE (5.3), of our pretrained MNMEF and fine-tuned MNMEF against four benchmark methods: EnKF, ESRF, LETKF, and IEnKF. We consider the Lorenz ’96 system with $\sigma_{y}=0.7$ (Figure 1(a)) and $\sigma_{y}=1.0$ (Figure 1(b)). In Figure 1(c), we show the relative improvement $\mathcal{E}_{\mathrm{RI}}$ (5.4) of our fine-tuned method compared to EnKF, IEnKF, and LETKF.
@@
Inflation and localization are very important for DA on the Lorenz ’96 system because of the high dimensionality of the states. All benchmark methods in this comparison utilized inflation, while only LETKF implemented localization. For the benchmark methods, the hyperparameters for both inflation and localization were optimized for different ensemble sizes, with the optimal values determined through grid search. Detailed results of this grid search can be found in B.3.
@@
We observe that our fine-tuned model consistently outperforms all benchmarks. Indeed our fine-tuned method achieves substantial improvement for small ensemble sizes and maintains a notable advantage at larger ensemble sizes, outperforming LETKF by approximately 15–20%. Our pretrained model, trained with $N=10$ ensembles also behaves competitively, but performs slightly worse than LETKF for large ensemble size $40,60,100$.
@@ 5.3 Kuramoto–Sivashinsky
The Kuramoto–-Sivashinsky (KS) equation [51, 52] is a widely studied nonlinear partial differential equation that describes the evolution of instabilities in spatially extended systems, capturing complex spatiotemporal chaotic dynamics arising in flame fronts, thin fluid films, and reaction-diffusion systems. Due to its chaotic behavior and sensitivity to initial conditions, the KS equation serves as a valuable testbed for developing and evaluating data assimilation algorithms. The one-dimensional KS equation for $u:[0,L]\times\mathbb{R}^{+}\to\mathbb{R}$ takes the form
$\frac{\partial u}{\partial t}+\frac{\partial^{4}u}{\partial x^{4}}+\frac{\partial^{2}u}{\partial x^{2}}+u\frac{\partial u}{\partial x}$ $=0,\quad(x,t)\in(0,L)\times(\infty),$ (5.6a)
$u(x,0)$ $=u_{0},\quad\forall x\in[0,L],$ (5.6b)
and periodicity in space is imposed to identity $x=L$ and $x=0.$ The detailed experimental settings and parameters for the KS system are summarized in Table 2. Numerically, the KS equation (5.6) is discretized spatially using a Fourier pseudo-spectral method to handle periodic boundary conditions effectively.
Table 2: Kuramoto–Sivashinsky (KS) System Settings
@@
Obs Time Step $\Delta t=1$
(a) $\sigma_{y}=0.7$
(b) $\sigma_{y}=1$
(c) Relative improvement
Figure 2: Comparison results on the Kuramoto–Sivashinsky (KS) system. Panels (a) and (b) show direct RMSE comparisons between different methods at observation noise levels $\sigma_{y}=0.7$ and $\sigma_{y}=1$, respectively. Panel (c) illustrates the relative improvement of our fine-tuning method compared to benchmarks. Our proposed MNMEF method is highlighted in the legends with bold font and an asterisk (e.g. Pretrain).
In Figure 2, we compare the performance of our pretrained model and fine-tuned model against four benchmark methods: EnKF, ESRF, LETKF, and IEnKF, for the KS system with $\sigma_{y}=0.7$ (Figure 2(a)) and $1.0$ (Figure 2(b)). Figure 2(c) illustrates the relative improvement of our fine-tuned method compared to EnKF, IEnKF, and LETKF.
Similarly to the Lorenz ’96 case (Subsection 5.2), inflation and localization play a crucial role in the KS system experiments. All benchmark methods employed inflation techniques, with LETKF being the only one with localization. Within these benchmark methods, the hyperparameters for both inflation and localization are optimized for each ensemble size (B.3).
@@
Remark 15 (Implement the Localization).
It is worth noting that methods other than LETKF could potentially accommodate localization techniques. For example, LETKF is essentially ESRF combined with localization. However, since our implementation is based on the DAPPER package, which only implements localization for LETKF, we do not consider localization for the other four methods. Our results on Lorenz ’96 and KS problems sufficiently demonstrate the advantages of our approach over other methods, even without localization. For larger ensemble sizes (e.g., $N=60,100$), where the benefits of localization diminish, our method still significantly outperforms the alternatives, like IEnKF.

### v2

@@ 5.1 Experimental Set-Up
In the following three Subsections 5.2, 5.3, and 5.4, we compare our method against several benchmark approaches:
@@
    EnKF Perturbed Observation (EnKF) [9, 10]: This is the classic implementation of EnKF that employs a stochastic update approach, where the observations are perturbed by adding random noise samples drawn from the observation noise distribution.
@@
    Ensemble Square Root Filter (ESRF) [11, 47]: This method is a deterministic variant of the EnKF. It directly transforms the ensemble using matrix square root operations to match the second order statistics of the EnKF, but without requiring stochastic perturbations.
@@
    Local Ensemble Transform Kalman Filter (LETKF) [14]: This method builds upon the ESRF by incorporating spatial localization. We employ the Gaspari–Cohn (GC) function [45] according to DAPPER package [48].
@@
    Pretrained MNMEF: Trained once at $N=10$ for $1000$ epochs and applied directly to all ensemble sizes, demonstrating transfer capability across different ensemble configurations. This appears as ‘Pretrain’ in our figures.
@@
    Fine-tuned MNMEF: Fine tune the pretrained model specific to each target ensemble size for $20$ epochs, as detailed in Subsection 4.4. This appears as ‘Tuned’ in our figures.
The intermediate feature dimensions in the set transformer for different dynamic models are provided in C.1. The detailed intiail conditions for different dynamic models are provided in C.3. The training hyperparameters including trajectory lengths, learning rates and batch sizes, are provided in C.4.
For fair comparison, each benchmark method is configured with optimized hyperparameters, which were determined through comprehensive grid search experiments conducted at various ensemble sizes. The detailed results of these grid search experiments, including parameter spaces explored and grid search heatmaps, are presented in C.5. This approach ensures that all benchmark methods achieve close to their optimal performance; the benchmarks thus present a robust test of our proposed methodology. Notably, we performed hyperparameter selection for the benchmarks directly on their test set performance, without using a validation set, which only strengthens the benchmark methods’ advantage in our comparisons. All benchmark results were generated using the DAPPER package [48].
The performance of the methods is evaluated using the difference between the estimated state (i.e., the mean of particles) and the ground truth state over the entire trajectory. In the continuous-time setting, let $\mathbf{v}^{\dagger}(t):[0,T]\rightarrow\mathbb{R}^{d_{v}}$ denote the ground truth trajectory between time 00 and $T$, and let $\mathbf{v}(t):[0,T]\rightarrow\mathbb{R}^{d_{v}}$ denote the estimated state trajectory. Let $\|\cdot\|_{2}$ denote the Euclidean norm in $\mathbb{R}^{d_{v}}$. Then, the performance can be evaluated using the relative $L^{1}:=L^{1}([0,T];\mathbb{R}^{d_{v}})$ error:
@@
In the discrete-time setting, we interpret the discrete time index $j$ as time $t_{j}=j\Delta t$, where $\Delta t=T/J$ is the time step size. The estimated trajectory is given by the sequence $V^{(N)}=\{\overline{v}^{(N)}_{j}\}_{j=1}^{J}$ where $\overline{v}^{(N)}_{j}=\frac{1}{N}\sum_{n=1}^{N}v^{(n)}_{j}$, and the ground truth trajectory is given by $V^{\dagger}=\{v_{j}^{\dagger}\}_{j=1}^{J}$. The relative discrete $L^{1}$ error can then be defined as
$\mathcal{E}(V^{(N)},V^{\dagger})=\frac{\sum_{j=1}^{J}\|\mathbf{v}(j\Delta t)-\mathbf{v}^{\dagger}(j\Delta t)\|_{2}\,\Delta t}{\sum_{j=1}^{J}\|\mathbf{v}^{\dagger}(j\Delta t)\|_{2}\,\Delta t}=\frac{\sum_{j=1}^{J}\|\overline{v}_{j}-v_{j}^{\dagger}\|_{2}}{\sum_{j=1}^{J}\|v_{j}^{\dagger}\|_{2}}.$ (5.2)
This error metric can be interpreted as a form of relative root-mean-square error (RMSE). At each time step $j$, the quantity $\|\overline{v}_{j}-v_{j}^{\dagger}\|_{2}$ corresponds to the root-mean-square error over the $d_{v}$-dimensional state vector. The overall metric thus represents a time-averaged spatial RMSE, normalized by the magnitude of the ground truth trajectory. In what follows, we refer to this error $\mathcal{E}(V^{(N)},V^{\dagger})$ as the relative RMSE (R-RMSE).
For $M_{\mathrm{test}}$ trajectories $\{V_{m}^{\dagger}\}_{m=1}^{M_{\mathrm{test}}}$ and their corresponding estimated trajectories $\{V_{m}^{(N)}\}_{m=1}^{M_{\mathrm{test}}}$, the averaged R-RMSE $\overline{\mathcal{E}}$ is given by:
¯$\overline{\mathcal{E}}=\frac{1}{M_{\mathrm{test}}}\sum_{m=1}^{M_{\mathrm{test}}}\mathcal{E}(V_{m}^{(N)},V_{m}^{\dagger}).$ (5.3)
In the experiments presented in the remainder of this section, we adopt the averaged R-RMSE $\overline{\mathcal{E}}$ as the primary evaluation metric, allowing us to ameliorate the effect of randomness inherent in each individual run and providing a more reliable assessment of the overall performance of different methods. Unless otherwise specified, the number of test trajectories is fixed at $M_{\mathrm{test}}=64$.
To further highlight the advantages of our method, we consider the relative improvement in terms of R-RMSE $\mathcal{E}_{\mathrm{RI}}$ defined as:
$\mathcal{E}_{\mathrm{RI}}=\frac{\overline{\mathcal{E}}_{\mathrm{benchmark}}-\overline{\mathcal{E}}_{\mathrm{ours}}}{\overline{\mathcal{E}}_{\mathrm{benchmark}}}.$ (5.4)
@@
We apply the following general hyperparameter setting in all of our experiments unless otherwise specified. The dynamic operator $\Psi$ in Equation (2.1) is obtained by solving a differential equation using the fourth-order Runge–Kutta (RK4) [49] method. The time step $\Delta t$ denotes the interval between consecutive observations. For accurate numerical integration of the dynamics, we perform multiple RK4 steps within each interval $\Delta t$, utilizing an integration step size smaller than $\Delta t$. The dynamic noise $\xi\sim\mathcal{N}(0,\Sigma)$ is typically set at $\Sigma=\sigma_{v}^{2}I$ where $I$ is the identity matrix and $\sigma_{v}=10^{-3}$ or $0$ For the observation noise, $\eta\sim\mathcal{N}(0,\Gamma)$, we define its covariance matrix as $\Gamma=\sigma_{y}^{2}I$, where we set $\sigma_{y}$ to either $0.7$ or $1.0$.
@@ 5.2 Lorenz ’96
$\frac{du_{i}}{dt}=\bigr{(}u_{i+1}-u_{i-2}\bigl{)}u_{i-1}-u_{i}+F,\quad i=1,2,\ldots,d.$ (5.5)
@@
States $v=(u_{1},u_{2},\ldots,u_{40})\in\mathbb{R}^{d},\quad d=40$
Observations $h(v)=(u_{1},u_{5},\cdots,u_{33},u_{37})\in\mathbb{R}^{d_{\mathrm{obs}}},\quad d_{\mathrm{obs}}=10$
Time Step Observation time step: $\Delta t=0.15$; $5$ RK4 integration steps: $\Delta t/5=0.03$
Relative RMSE Refer to caption Refer to caption
Relative Improvement Refer to caption Refer to caption
Lorenz ’96 Ensemble Size Ensemble Size
$\sigma_{y}=0.7$ $\sigma_{y}=1.0$
@@
Figure 1: Comparison results on the Lorenz ’96 system. The upper row of plots shows direct R-RMSE comparisons between different methods, while the lower row illustrates the relative improvement of our fine-tuning method compared to benchmarks. These comparisons are presented for observation noise levels $\sigma_{y}=0.7$ (left column) and $\sigma_{y}=1.0$ (right column). Our proposed MNMEF method is highlighted in the legends with bold font and an asterisk (e.g. Pretrain). In summary, our fine-tuned MNMEF model consistently outperforms benchmarks, showing substantial improvements for small ensembles and a 15-20% advantage over LETKF at larger ensemble sizes (eg. 60, 100).
Ground Truth Refer to caption Observe every 4th dimension; $\sigma_{y}=1.0$${\hskip 85.35826pt}$ Refer to caption
(a) States (b) Observations
MNMEF (ours) $N=10$ Refer to caption Refer to caption Refer to caption
LETKF [14] $N=10$ Refer to caption Refer to caption Refer to caption
(c) States (mean) (d) Abs Error (e) Spread (std)
Figure 2: Visualization of one test trajectory (time steps 1401-1500, $\Delta t=0.15$) for Lorenz ’96 states (vertical axis) over time (horizontal axis)with the observation noise $\sigma_{y}=1.0$. Panel (a): Ground-truth states (unknown). Panel (b): Observations (known, every 4th dimension observed). Rows 2-3 show our method MNMEF pretrained on $N=10$, and the benchmark LETKF with the ensemble size $N=10$. Panel (c): state estimation (ensemble mean), Panel (d): absolute error of mean with respect to the ground truth, and Panel (e): ensemble spread (standard deviation).
MNMEF (ours) $N=10$ Refer to caption Refer to caption
LETKF [14] $N=10$ Refer to caption Refer to caption
(a) Dimension 1 (Observed) (b) Dimension 2 (Not Observed)
Figure 3: Visualization of two dimensions (index 1 and 2) in one test trajectory (time steps 1401-1500, $\Delta t=0.15$) for Lorenz ’96 state values (vertical axis) over time (horizontal axis) with the ensemble size $N=10$ and the observation noise $\sigma_{y}=1.0$. Panel (a): Dimension 1, observed. Panel (b): Dimension 2, not observed. The first row is our method MNMEF, pretrained on $N=10$, and the second row is the benchmark LETKF. In summary, MNMEF performs significantly better on the unobserved dimension and comparably to LETKF on the observed dimension; meanwhile, MNMEF maintains an appropriate spread without suffering from filter degeneracy.
@@
In Figure 1, we compare the performance of our pretrained MNMEF and fine-tuned MNMEF against four benchmark methods: EnKF, ESRF, LETKF, and IEnKF, on the Lorenz ’96 system. The upper row of plots in Figure 1 displays the R-RMSE (5.3) for observation noise levels $\sigma_{y}=0.7$ (left plot) and $\sigma_{y}=1.0$ (right plot). The lower row of plots in Figure 1 shows the relative improvement $\mathcal{E}_{\mathrm{RI}}$ (5.4) of our fine-tuned method as compared to EnKF, IEnKF, and LETKF for these respective noise levels.
@@
Inflation and localization are important for DA on the Lorenz ’96 system because of the high dimensionality of the states. All benchmark methods in this comparison utilized inflation, while only LETKF implemented localization. For the benchmark methods, the hyperparameters for both inflation and localization were optimized for different ensemble sizes, with the optimal values determined through grid search. Detailed results of this grid search can be found in C.5.
@@
We observe that our fine-tuned model consistently outperforms all benchmarks. Indeed, our fine-tuned method achieves substantial improvement for small ensemble sizes and maintains a notable advantage at larger ensemble sizes, outperforming LETKF by approximately 15–20%. Our pretrained model, trained with $N=10$ ensembles also behaves competitively, but performs slightly worse than LETKF for large ensemble size $40,60,100$.
To further illustrate the performance differences between our MNMEF and the best-performance benchmark LETKF, we provide visualizations in Figure 2 and Figure 3 for $N=10$ and $\sigma_{y}=1.0$. The visualization is on the last $100$ time steps (from $1401$ to $1500$) from a test trajectory of length $1500$ and $\Delta t=0.15$. In both figures, our MNMEF approach is the one pretrained with the ensemble size $N=10$.
Figure 2 presents a comparison of the estimated state trajectory. Panel (a) shows the ground truth, and panel (b) shows the sparse observations. The subsequent rows compare MNMEF and LETKF. Visually, MNMEF displays smaller magnitudes for the absolute error (panel (d)). In addition, the ensemble spread (panel (e)) for MNMEF is more consistent than LETKF. Figure 3 focuses on the estimation of two specific dimensions: one observed (Dimension 1, panel (a)) and one unobserved (Dimension 2, panel (b)). For the observed dimension, both MNMEF and LETKF perform comparably well, with their ensemble means closely following the ground truth. However, for the unobserved dimension, MNMEF demonstrates a clear advantage in state estimation. Furthermore, MNMEF maintains an appropriate ensemble spread while avoiding filter degeneracy.
@@ 5.3 Kuramoto–Sivashinsky
The Kuramoto–-Sivashinsky (KS) equation [51, 52] is a widely studied nonlinear partial differential equation that describes the evolution of instabilities in spatially extended systems, capturing complex spatiotemporal chaotic dynamics arising in flame fronts, thin fluid films, and reaction-diffusion systems. Due to its chaotic behavior and sensitivity to initial conditions, the KS equation serves as a valuable testbed for developing and evaluating data assimilation algorithms. With the periodicity in space imposed to identity $x=L$ and $x=0$, the one-dimensional KS equation for $u:[0,L]\times\mathbb{R}^{+}\to\mathbb{R}$ takes the form
$\frac{\partial u}{\partial t}+\frac{\partial^{4}u}{\partial x^{4}}+\frac{\partial^{2}u}{\partial x^{2}}+u\frac{\partial u}{\partial x}$ $=0,\quad(x,t)\in(0,L)\times\mathbb{R}^{+},$ (5.6a)
$u(x,0)$ $=u_{0},\quad\forall x\in[0,L].$ (5.6b)
Table 2: Kuramoto-–Sivashinsky (KS) System Settings
@@
Time Step Observation time step: $\Delta t=1$; $4$ RK4 integration steps: $\Delta t/4=0.25$
Relative RMSE Refer to caption Refer to caption
Relative Improvement Refer to caption Refer to caption
Kuramoto– Ensemble Size Ensemble Size
Sivashinsky $\sigma_{y}=0.7$ $\sigma_{y}=1.0$
Figure 4: Comparison results on the Kuramoto–-Sivashinsky (KS) system. The upper row of plots shows direct R-RMSE comparisons between different methods, while the lower row illustrates the relative improvement of our fine-tuning method compared to benchmarks. These comparisons are presented for observation noise levels $\sigma_{y}=0.7$ (left column) and $\sigma_{y}=1.0$ (right column). Our proposed MNMEF method is highlighted in the legends with bold font and an asterisk (e.g. Pretrain). In summary, our fine-tuned MNMEF model consistently outperforms benchmarks, showing substantial improvements for small ensembles and around 20% advantage over LETKF at larger sizes (eg. 60, 100).
Ground Truth Refer to caption Observe every 8th dimension; $\sigma_{y}=1.0$${\hskip 85.35826pt}$ Refer to caption
(a) States (b) Observations
MNMEF (ours) $N=10$ Refer to caption Refer to caption Refer to caption
LETKF [14] $N=10$ Refer to caption Refer to caption Refer to caption
(c) States (mean) (d) Abs Error (e) Spread (std)
Figure 5: Visualization of one test trajectory (time steps 1901-2000, $\Delta t=1$) for Kuramoto–Sivashinshy (KS) states (vertical axis) over time (horizontal axis) with the observation noise $\sigma_{y}=1.0$. Panel (a): Ground-truth states (unknown). Panel (b): Observations (known, every 8th dimension observed). Rows 2-3 show our method MNMEF pretrained on $N=10$, and the benchmark LETKF with the ensemble size $N=10$. Panel (c): state estimation (ensemble mean), Panel (d): absolute error with ground truth, and Panel (e): ensemble spread (standard deviation).
MNMEF (ours) $N=10$ Refer to caption Refer to caption
LETKF [14] $N=10$ Refer to caption Refer to caption
(a) Dimension 1 (Observed) (b) Dimension 2 (Not Observed)
Figure 6: Visualization of two dimensions (index 1 and 2) in one test trajectory (time steps 1901-2000, $\Delta t=1$) for Kuramoto–Sivashinshy (KS) state values (vertical axis) over time (horizontal axis) with the ensemble size $N=10$ and the observation noise $\sigma_{y}=1.0$. Panel (a): Dimension 1, observed. Panel (b): Dimension 2, not observed. The first row is our method MNMEF, pretrained on $N=10$, and the second row is the benchmark LETKF. In summary, MNMEF performs slightly better on both the observed and the unobserved dimensions; meanwhile, MNMEF maintains an appropriate spread without suffering from filter degeneracy.
The detailed experimental settings and parameters for the KS system are summarized in Table 2. Numerically, the KS equation (5.6) is discretized spatially using a Fourier pseudo-spectral method to handle periodic boundary conditions effectively.
In Figure 4, we compare the performance of our pretrained model and fine-tuned model against four benchmark methods: EnKF, ESRF, LETKF, and IEnKF, for the KS system with $\sigma_{y}=0.7$ and $1.0$, and provide the relative improvement of our fine-tuned method as compared to EnKF, IEnKF, and LETKF.
Similarly to the Lorenz ’96 case (Subsection 5.2), inflation and localization play a crucial role in the KS system experiments. All benchmark methods employed inflation techniques, with LETKF being the only one with localization. Within these benchmark methods, the hyperparameters for both inflation and localization are optimized for each ensemble size (C.5).
@@
Similarly to the Lorenz ’96 discussion in Subsection 5.2, we provide visualizations in Figure 2 and Figure 3 for $N=10$ and $\sigma_{y}=1.0$ to further illustrate the performance differences between our MNMEF and the best-performance benchmark LETKF on the KS system. The visualization is on the last $100$ time steps (from $1901$ to $2000$) of a test trajectory with length $2000$ and $\Delta t=1$. In both figures, our MNMEF is the one pretrained with the ensemble size $N=10$.
Figure 5 presents a comparison of the estimated state trajectory. Figure 6 visualizes the estimation of two specific dimensions: one observed (Dimension 1, panel (a)) and one unobserved (Dimension 2, panel (b)). In summary, our MNMEF performs slightly better than LETKF on both the observed and the unobserved dimensions, which is verified by the R-RMSE quantitative comparison in Figure 4. MNMEF maintains an appropriate ensemble spread, similarly to the LETKF.
Remark 16 (Implementation of Localization).
It is worth noting that methods other than LETKF could potentially accommodate localization techniques. For example, LETKF is essentially the ESRF combined with localization. However, since our implementation is based on the DAPPER package, which only implements localization for LETKF, we do not consider localization for the other four methods. Our results on Lorenz ’96 and KS problems sufficiently demonstrate the advantages of our approach over other methods, even without localization. For larger ensemble sizes (e.g., $N=60,100$), where the benefits of localization diminish, our method still significantly outperforms the alternatives, like IEnKF.